A business invests $10,000 in a savings account for two years. At the beginning of the second year, an additional $3500 is invested. At the end of the second year, the account balance is $15,569.75. What was the annual interest rate?
Answer by checkley71(8403)

10000+(10000*2X)+3500+(3500*X)=15569.75 OR 13500+20000X+3500X=15569.75 OR
23500X=15569.75-13500 OR 23500X=2069.75 OR X=2069.75/23500 OR X=.0880744 OR THE INTEREST IS 8.8% PER ANNUM.
PROOF 10000*1.1761488+3500*1.0880744=15569.75 OR 11761.488+3808.260=15569.75